A 36.165 mg sample of a chemical known to contain only carbon, hydrogen, sulfur, and oxygen is put into a combustion analysis apparatus, yielding 64.425 mg of carbon dioxide and 26.373 mg of water. In another experiment, 35.809 mg of the compound is reacted with excess oxygen to produce 15.47 mg of sulfur dioxide.

Add subscripts to the formula provided to correctly identify the empirical formula of this compound. Do not change the order of the elements.
Answer #1

Giyer - 36.165 mg (C,H, S, og combustion → Co2 64.425 + H₂O 26.373 my mg 36 roy my combustion - excess S02 1547 mg solution ③

2do composition do of c in 30.1K5 my sample = 1756 x 100 = 48.55 do 36. 165 8.16 do do of H in 36-165 my somple = 295 x 100 =3BS Divide by smallest whte number c - 4.042/ 0.614 = 5.997 260 H - 8.09 / 0-674 = 120 s- 0.676/0.874 =1 boot o- 1.36%.874 = 2